### v2

# Continual Vision-and-Language Navigation

Abstract
In developing Vision-and-Language Navigation (VLN) agents that navigate to a destination using natural language instructions and visual cues, current studies largely assume a train-once-deploy-once strategy. We argue that this kind of strategy is less realistic, as deployed VLN agents are expected to encounter novel environments continuously through their lifetime. To facilitate more realistic setting for VLN agents, we propose Continual Vision-and-Language Navigation (CVLN) paradigm for agents to continually learn and adapt to changing environments. In CVLN, the agents are trained and evaluated incrementally across multiple scene domains (i.e., environments). We present two CVLN learning setups to consider diverse forms of natural language instructions: Initial-instruction based CVLN, focused on navigation via initial-instruction interpretation, and dialogue-based CVLN, designed for navigation through dialogue with other agents. We introduce two simple yet effective baseline methods, tailored to the sequential decision-making needs of CVLN: Perplexity Replay (PerpR) and Episodic Self-Replay (ESR), both employing a rehearsal mechanism. PerpR selects replay episodes based on episode difficulty, while ESR stores and revisits action logits from individual episode steps during training to refine learning. Experimental results indicate that while existing continual learning methods are insufficient for CVLN, PerpR and ESR outperform the comparison methods by effectively utilizing replay memory.
1 Introduction
One of the desired attributes of robotic agents is their ability to learn and adapt to new environments without losing knowledge of previous ones. In practice, robots deployed in the human world are required to perform a variety of tasks in complex and changing environments. Vision-and-Language Navigation (VLN) focuses on agents that navigate using natural language instructions and visual cues, promising for applications like robotic assistants. Despite recent advancements, most VLN research follows a train-once-deploy-once strategy.
In real-world applications, environments are not static; they evolve over time with new buildings, renovations, or changes in layout. A deployed VLN agent must promptly adapt whenever the environment changes throughout its lifetime. Finetuning on new environments might seem practical, but it can cause the agent’s performance in previously learned contexts to deteriorate, a challenge known as catastrophic forgetting [30]. To address this challenge, the agent needs the capability for Continual Learning (CL) [42], which enables it to learn and adapt to new circumstances while retaining its existing knowledge.
In this paper, we propose Continual Vision-and-Language Navigation (CVLN) paradigm, which trains and evaluates VLN agents with consideration of the CL ability. In CVLN, agents explore different scene domains—each comprising various indoor scenes with multiple navigation episodes—sequentially and are evaluated on their ability to navigate through all of them. The CVLN agents need to retain their ability on previously learned scene domains while learning the current scene domain. Figure 2 illustrates an overview of CVLN. In CVLN, the agent incrementally learns multiple scene domains, and as shown in Figure 1, there are substantial differences in visual distributions among scene domains. These distributional discrepancies can lead to catastrophic forgetting when the domains are learned incrementally.
The CVLN supports two settings to handle different types of natural language instructions: Initial-instruction based CVLN (I-CVLN) and Dialogue based CVLN (D-CVLN). In I-CVLN, the instructions given to the agent at the start of navigation contain fine-grained information about the entire navigation path. In D-CVLN, the agent receives a target object and a conversation history between humans collaborating to locate it, and must reason navigation actions to reach the goal within environments.
We present two simple yet effective baseline methods for CVLN—Perplexity Replay (PerpR) and Episodic Self-Replay (ESR)—both based on a rehearsal mechanism, where the effective construction and utilization of replay memory are crucial. PerpR uses self-evaluated action perplexity (i.e. difficulty) to assess the importance of episodes, effectively selecting and organizing them in the replay memory based on their action perplexity. After learning a scene domain, the agent adds high-perplexity episodes to the replay memory and removes low-perplexity ones. Meanwhile ESR stores action logits predicted by the agent for each individual step within episodes in the replay memory. During subsequent training, the agent revisits these stored logits, allowing it to refine its learning by leveraging detailed episode step-level logits from past experiences.
To validate the effectiveness of our methods, we conducted comparative experiments with existing CL methodologies. Furthermore, we conducted an ablation study, analyzed the effects of varying replay memory sizes, and examined the stability-plasticity trade-off.
Figure 1: Frechet Inception Distance[16] heatmaps illustrating the distribution shifts between visual features [47] across 10 scene domains for the I-CVLN and D-CVLN.
To summarize our contributions:
  • We propose a Continual Vision-and-Language Navigation (CVLN) paradigm to develop VLN agents capable of continual learning, introducing two settings— I-CVLN (instruction-based) and D-CVLN (dialogue-based)—for comprehensive evaluation across diverse natural language instruction types.
  • We propose two simple yet effective CVLN methods, PerpR and ESR, which enhance continual learning through rehearsal mechanisms, applicable to tasks like CVLN that involve sequential decision-making.
  • In experiments, PerpR and ESR demonstrate their effectiveness in the CVLN paradigm, achieving improved performance over comparative methods in both I-CVLN and D-CVLN settings.
Figure 2: Comparison between (a) VLN and (b) CVLN. In VLN, the agent is trained within fixed scene domains and then evaluated on unseen scene domains. In contrast, the agent in CVLN is trained within a new scene domain sequentially and and evaluated on those scene domains along with previously encountered ones.
2 Related Work
Continual Learning. Continual Learning (CL) focuses on training models to acquire new knowledge over time while retaining previously learned information. CL methods have generally been extensively studied for simple tasks [13] in computer vision. This field includes several CL scenarios. Instance-Incremental Learning trains on batches of samples for the same task [27]. Domain-Incremental Learning (DIL) deals with tasks having the same labels but different input distributions, with no domain identity during inference [19]. Task-Incremental and Class-Incremental Learning involve different label spaces, with varying use of domain identity [19]. Task-Free Continual Learning handles separate label spaces without domain identity [1], and Online Continual Learning processes sequential data streams in real-time [2]. CVLN fits within DIL, providing scene domain identities during training but not during testing, requiring generalization without explicit domain knowledge at evaluation. There are three main approaches for CL: Regularization [22], which adds terms to prevent the loss of prior knowledge; Rehearsal [34, 9, 8, 7], which stores and reuses data from past tasks; and Architectural [36, 29], which introduces task-specific parameters to the agent’s architecture. This paper proposes two Rehearsal-based methods, PerpR and ESR, and evaluates their effectiveness through comparative experiments.
Vision-and-Language Navigation. In VLN, agents navigate realistic 3D indoor environments by following natural language instructions. Various datasets like R2R [6], R4R [20], and RxR [24] have been developed, focusing on navigation based on initial instructions that fully describe the path. Other tasks, such as CVDN [41] and HANNA [32], involve agents obtaining navigation information through dialogue with an oracle or other agents. Additionally, datasets like Touchdown [10] and StreetLearn [31] have expanded VLN research to outdoor environments. IVLN [23] present a paradigm where agents maintain memory across episodes to navigate a sequence of episodes. In the field of VLN, research focuses on improving agent architecture [17, 11, 14, 4], enhancing visual-text alignment through auxiliary losses [28, 46, 48], using data augmentation [40, 45, 25, 21], and extensive pre-training [15, 11, 12, 3] to enhance generalization. In this paper, we adopted VLN-BERT [17] and HAMT [11] as the backbone architectures. Existing VLN datasets and methods lack considerations for CL, highlighting the need for scalable strategies that enable agents to learn efficiently to diverse, changing environments without costly retraining models from scratch.
3 Continual Vision-and-Lauguage Navigation
3.1 Formulation of VLN
The existing VLN agents are trained on fixed training data and evaluated on episodes from scenes that were not encountered during training. The objective of VLN agents is to minimize the following loss:
$\mathcal{L}_{D}\triangleq\mathbb{E}_{(I,A^{*})\sim D}[\ell(\pi_{\theta}(V,I),A^{*})]$ (1)
where $\theta$ specifies the parameters for the VLN agent, $\pi$. The dataset $D$ is composed of episodes, which include navigation instructions $I$ and their corresponding ground-truth trajectory $A^{*}=\{a_{1}^{*},a_{2}^{*},\ldots,a_{N}^{*}\}$. The set $V=\{v_{1},v_{2},\ldots,v_{N}\}$ represents the sequence of visual observations from the environment, each associated with a step determined by a predicted action, with $N$ denoting the total number of steps per episode. The agent $\pi_{\theta}(V,I)$ determines the actions based on $V$ and $I$ sequentially. The navigation imitation loss function $\ell$ then compares this behavior to the ground-truth trajectory $A^{*}$ to evaluate the discrepancy between them. We use cross entropy loss for $\ell$.
3.2 Formulation of CVLN
A scene domain includes episodes in multiple indoor environments, and each episode comprises a natural language navigation instruction and its corresponding trajectory. CVLN agent learns scene domains incrementally. The agent is evaluated on the incrementally learned scene domains. The evaluation episodes belong to the learned scene domains but are episodes that have not been seen during training. During this process, the agent must learn the current scene domain without forgetting the knowledge about previously learned scene domains. This learning approach is more suitable for real-world agents that encounter and need to learn new scene domains continually.
Formally, a CVLN is divided into $S$ scene domains; during each scene domain $s\in\{1,\ldots,S\}$, an episode $\epsilon$, navigation instruction $I$ and their corresponding ground-truth trajectory $A^{*}$ are drawn from an independent and identically distributed (i.i.d.) distribution $D_{s}$. The objective of agents in CVLN is as follows:
[…]
$\quad\text{where}\ \mathcal{L}_{D_{s}}\triangleq\mathbb{E}_{(I,A^{*})\sim D_{s}}[\ell(\pi_{\theta}(V,I),A^{*})]$
Figure 3: Comparison between I-CVLN and D-CVLN. In I-CVLN, the agent is given an initial instruction containing all the information about the navigation path. Conversely, in D-CVLN, the agent obtains information about the navigation path through communication with an oracle.
where an agent $\pi$, with parameters $\theta$, is optimized on one scene domain at a time in a sequential manner. $s_{c}$ represents the current scene domain.
3.3 Initial-Instruction based CVLN
In I-CVLN, agents receive initial instructions outlining the entire navigation path. As the R2R [6] and RxR [24] datasets provide initial instruction-based navigation episodes, we construct our continual learning benchmark for I-CVLN on top of these datasets, using only english epidodes. The training split from the original datasets is used for I-CVLN training, while the val-seen split, which contain episodes from the same scenes but not seen during training, is used for I-CVLN validation. We grouped episodes from the same scenes across both datasets into 20 scene domains, each containing over 600 episodes for training.
Datasets Split Episode Scene Domain
I-CVLN Train 15700 20
Valid 1563 20
D-CVLN Train 3737 11
Valid 382 11
Table 1: Dataset statistics for I-CVLN and D-CVLN. The table outlines the number of episodes and scene domains available in the training (Train) and validation (Valid) splits for each dataset.
3.4 Dialogue based CVLN
In D-CVLN, agents navigate from dialogue history. As the CVDN [41] dataset includes dialogue data between a follower and an oracle paired with corresponding trajectories, we build benchmark for D-CVLN on top of it. Due to the limited episodes per scene in CVDN, we grouped four scenes into each scene domain, resulting in 11 domains for training. These scene domains are learned sequentially. The training split of CVDN is used for D-CVLN training, while the val-seen split, containing evaluation episodes for each domain but not seen during training, is used for D-CVLN validation. Figure 3 illustrates a comparison between I-CVLN and D-CVLN.
3.5 Evaluation Protocol for CVLN
The evaluation of CVLN agents focuses on assessing their ability to retain previously learned knowledge while effectively adapting to new scene domain. For CVLN evaluation, the Average Metric is used to evaluate agent performance across different scene domains. The average metric is defined as follows:
$\frac{1}{S}\sum_{i=1}^{S}R_{S,i}$ (3)
where $R_{S,i}$ indicates the agent’s performance on the $i$th domain after training the $S$th scene domain. The metric evaluates the agent’s knowledge retention across prior scene domains and its adaptability to new ones.
For I-CVLN, we use standard VLN metrics [5] $R_{S,i}$ in Equation 3: Success Rate (SR), Success weighted by inverse Path Length (SPL), and Navigation Error (NE). For D-CVLN, we use the main metric from CVDN [41], $R_{S,i}$ in Equation 3, which is Goal Progress (GP) in meters.
4 Methods
Rehearsal-based CL approaches mitigate catastrophic forgetting by using a fixed-size replay memory $M$ to store and revisit episodes from previous domains. After training each task, $M$ is updated by deleting an equal number of old episodes and adding new ones from the latest domain. Effective replay memory construction and usage are crucial for maintaining performance across domains. In this context, we propose Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) as baseline methods for the CVLN agent. Figure 4 provides an overview of PerpR and ESR.
Figure 4: Overview of Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) for CVLN agents. (a) PerpR prioritizes challenging episodes in the agent’s memory, optimizing for high Action Perplexity (AP). (b) ESR enables the agent to self-replay using past optimal behaviors.
4.1 Perplexity Replay
To construct replay memory for continual learning, random sampling [44, 35] is often employed. However, this approach fails to consider the importance of samples in recovering past knowledge, resulting in suboptimal replay memory composition. Existing methods [39, 2] have attempted to address this by utilizing gradient-based or optimizable approaches to construct more effective replay memory. Nevertheless, these methods are typically tailored to simple tasks such as image classification and are not suitable for tasks that require sequential action prediction, such as CVLN.
Algorithm 1 Perplexity Replay Memory Update
1:  Input: Replay memory $M$, replay memory size $\lvert M\rvert$, set of episodes from the last scene domain $E_{\text{last}}$, agent parameters $\theta$, the number of trained scene domains $s$
2:  Output: Updated replay memory $M$
3:  Calculate Action Perplexity (AP) for each episode in $E_{\text{last}}$ and $M$
[…]
6:  $n_{s}\leftarrow\lvert M\rvert/(s+1)$
7:  for $s=1$ to $s$ do
8:     Remove episodes with the lowest AP from $M$ for domain $s$ until only $n_{s}$ episodes remain
9:  end for
10:  Add episodes from $E_{\text{last}}$ to $M$ until $\lvert M\rvert$ equals fixed size
11:  return  $M$
To overcome these limitations, we propose Perplexity Replay (PerpR), which constructs the replay memory by quantifying the difficulty of VLN episodes. A key innovation of PerpR is that the model self-evaluates the difficulty of each episode using Action Perplexity (AP), without relying on external metrics or annotations. By self-assessing its own uncertainty through AP, the model identifies challenging episodes where it is less confident. PerpR then prioritizes these difficult episodes in the replay memory, enabling the agent to focus on experiences that require further learning and thus enhancing the overall learning process.
Assuming episodes with high AP indicate insufficient agent learning, PerpR ensures that the replay memory is populated with experiences most beneficial for the agent’s continued improvement. The replay memory update procedure involves selecting episodes with low AP for removal and adding those with high AP from previously trained scene domains. This self-directed mechanism allows the agent to concentrate on challenging episodes it has identified post-training, thereby improving its performance. The Action Perplexity is computed as:
$\text{AP}_{\theta}(\epsilon)=\exp\left({-\frac{1}{N}\sum_{i=1}^{N}\log P_{\theta}(a_{i}|v_{i},I)}\right)$ (4)
where $N$ is the number of steps in an episode, $v_{i}$ is the observed visual state at step $i$, and $P(a_{i}|v_{i},I,\theta)$ is the probability of predicted action $a_{i}$ of agent $\pi$ parameterized by $\theta$ based on $v_{i}$, the instructions $I$.
We combine the replay memory with data from the current scene domain for the training of subsequent scene domains. The loss to be minimized is formulated as follows:
$\mathcal{L}_{PerpR}=\mathcal{L}_{D_{s_{c}}}+\mathcal{L}_{M}$ (5)
$\mathcal{L}_{D_{s_{c}}}$ $\triangleq\mathbb{E}_{(I,A^{*})\sim D_{s_{c}}}[\ell(\pi_{\theta}(V,I),A^{*})]$
$\mathcal{L}_{M}$ $\triangleq\mathbb{E}_{(I,A^{*})\sim M}[\ell(\pi_{\theta}(V,I),A^{*})]$
where $\mathcal{L}_{D_{s_{c}}}+\mathcal{L}_{M}$ represents the sum of the losses over the current scene domain $D_{s_{c}}$ and the replay memory $M$. The memory update procedure for PerpR is described in Algorithm 1.
4.2 Episodic Self-Replay
To effectively utilize the samples in the replay memory, prior works [34, 8, 7, 43] have employed logit distillation methods, which were designed for continual learning in non-sequential tasks such as image classification or object detection. However, they are not directly applicable to CVLN due to its sequential action prediction in dynamic environments. In CVLN, agents must perform sequences of actions where observations depend on previous actions, involving temporal state changes and decision-making processes. Existing logit distillation methods focus on single-step predictions, limiting their effectiveness in learning or generalizing complex sequential action patterns.
To address this limitation, we extended these approaches for application to CVLN episodes. At each action step, we stored the action logits in a replay memory. When learning from episodes in the replay memory, the CVLN agent utilizes these stored logits to train itself to follow its past action distribution at each step. To reinforce consistency, we applied student forcing [38] during training by using the stored logits as the next action for subsequent steps, rather than the predicted logits. This encourages adherence to prior action patterns and aligns the learning process with the sequential nature of CVLN.
Algorithm 2 Episodic Self-Replay
0:  Dataset $D_{s}$, Memory $M$, Learning rate $\eta$, Agent $f_{\theta}$, $\text{GetObservation}(t)$, $\text{ActionExecuter}(a_{t})$, $\lambda_{1}$, $\lambda_{2}$
1:  Sample $(I,A^{*})\sim D_{s}$
2:  $\mathcal{L}_{D_{c}}\leftarrow 0$, $Z\leftarrow[]$, $t\leftarrow 1$
3:  $v_{t}\leftarrow\text{GetObservation}(t)$;
4:  while $a_{t}$ does not predict stop do
5:     $z_{t}\leftarrow f_{\theta}(I,v_{t})$; $Z.\text{append}(z_{t})$
6:     $a_{t}\leftarrow\arg\max(z_{t})$
7:     $a_{t}^{*}\leftarrow A^{*}(t)$ //Get true action
8:     $\mathcal{L}_{D_{c}}\leftarrow\mathcal{L}_{D_{c}}+\ell(a_{t},a_{t}^{*})$
9:     $v_{t+1}\leftarrow\text{ActionExecuter}(a_{t})$
10:     $t\leftarrow t+1$
11:  end while
12:  $\mathcal{L}_{D_{c}}\leftarrow\mathcal{L}_{D_{c}}/t$
13:  Update $M\leftarrow\text{Random}(M,(I,A^{*},Z))$
14:  Sample $(I^{\prime},A^{\prime*},Z^{\prime*})\sim M$
15:  $\mathcal{L}_{ESR}\leftarrow 0$, $\mathcal{L}_{M}\leftarrow 0$, $t\leftarrow 1$
16:  $v_{t}\leftarrow\text{GetObservation}(t)$;
17:  while $a_{t}$ does not predict stop do
18:     $z_{t}\leftarrow f_{\theta}(I^{\prime},v_{t})$
19:     $a_{t}\leftarrow\arg\max(z_{t})$
20:     $a_{t}^{\prime*}\leftarrow A^{\prime*}(t)$ //Get true action
21:     $\mathcal{L}_{M}\leftarrow\mathcal{L}_{M}+\ell(a_{t},a_{t}^{\prime*})$
22:     $z_{t}^{\prime*}\leftarrow Z^{\prime*}(t)$ //Get replay logit
23:     $\mathcal{L}_{ESR}\leftarrow\mathcal{L}_{ESR}+\|z_{t}-z_{t}^{\prime*}\|^{2}_{F}$//Logit distillation
24:     $a_{t}^{*}\leftarrow\arg\max(z_{t}^{\prime*})$ //Get replay logit’s action
25:     $v_{t+1}\leftarrow\text{ActionExecuter}(a_{t}^{*})$
26:     $t\leftarrow t+1$
27:  end while
28:  $\mathcal{L}_{ESR}\leftarrow\mathcal{L}_{ESR}/t$
29:  $\mathcal{L}_{M}\leftarrow\mathcal{L}_{M}/t$
30:  $\mathcal{L}_{ESR\_{total}}\leftarrow\mathcal{L}_{D_{c}}+\lambda_{1}\mathcal{L}_{ESR}+\lambda_{2}\mathcal{L}_{M}$
31:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{ESR\_{total}}$
We define the agent parameters immediately after learning the scene domain $s$ as $\theta_{s}^{*}$. A function $f$ outputs the logits for action candidates at each sequential step of navigation. Unlike other rehearsal-based methods, ESR stores the sequential action logits $Z\triangleq\{z_{n}\}_{n=1}^{N}\triangleq f_{\theta_{s}^{*}}(V,I)$, corresponding to each step in the sequence, when updating the replay memory $M$. ESR’s replay memory uses random sampling [44] to select episodes to add and delete.
The ESR loss function can be formulated as follows:
$\mathcal{L}_{ESR}\triangleq\mathbb{E}_{(I,Z)\sim M}\left[\sum_{n=1}^{N}\|z_{n}-f_{\theta}(V_{n},I_{n})\|^{2}_{F}\right]$ (6)
Method I-CVLN D-CVLN
[…]
1. Vanilla $23.1\pm 0.8$ $25.6\pm 0.6$ $10.5\pm 0.5$ $5.5\pm 0.4$
2. Joint $40.1\pm 0.2$ $42.6\pm 0.1$ $7.1\pm 0.3$ $8.1\pm 0.1$
3. L2 $13.3\pm 0.2$ $14.6\pm 0.2$ $12.7\pm 0.3$ $4.3\pm 0.4$
4. AdapterCL $7.8\pm 0.3$ $9.1\pm 0.6$ $13.9\pm 0.8$ $4.3\pm 1.2$
5. AGEM $4.6\pm 2.8$ $4.5\pm 2.2$ $13.0\pm 0.2$ $2.7\pm 0.9$
6. RandR $24.9\pm 0.6$ $28.0\pm 0.7$ $10.0\pm 0.5$ $5.7\pm 0.3$
7. PerpR (Ours) $\underline{26.1}\pm 1.6$ $\underline{28.9}\pm 1.8$ $\underline{9.8}\pm 0.2$ $\textbf{6.1}\pm 0.4$
8. ESR (Ours) $\textbf{28.2}\pm 0.4$ $\textbf{31.9}\pm 0.9$ $\textbf{9.2}\pm 0.3$ $\underline{5.8}\pm 0.5$
Table 2: Comparative results of various methods on I-CVLN and D-CVLN. This table reports the mean and standard deviation values across three distinct learning curricula.
ESR uses $\mathcal{L}_{ESR}$ to allow agents to replay their behavior distribution during the learning process, which allows them to learn about new scene domains while effectively retaining previously learned knowledge. This helps mitigate catastrophic forgetting and improves the agent’s ability to learn and adapt efficiently in a constantly changing environment. Also, ESR employs $\mathcal{L}_{M}$, utilizing ground truth actions from replay memory. The final loss function of ESR is defined as follows:
$\mathcal{L}_{ESR\_total}=\mathcal{L}_{D_{s_{c}}}+\lambda_{1}\mathcal{L}_{M}+\lambda_{2}\mathcal{L}_{ESR}$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ control the impact of $\mathcal{L}_{M}$ and $\mathcal{L}_{ESR}$ on the overall loss. Both $\lambda_{1}$ and $\lambda_{2}$ are set to 0.2. Please refer to Algorithm 2 for more details.
5 Experiments
5.1 Comparative Methods
To establish the upper limit of our results, we provide the results of Joint agent trained across all scene domains simultaneously. Conversely, for the lower limit, we provide the result of Vanilla agent trained on scene domains sequentially without incorporating any CL methods.
L2 Regularization reduces the difference between the current parameter $\theta_{s}$ and the previous $\theta_{s-1}$ by adding the regularization term $\lambda\lVert\theta_{s}-\theta_{s-1}^{*}\rVert^{2}$ to the loss function.
Random Replay (RandR) uses Reservoir sampling [44] to manage replay memory, integrating selected episodes with current training scene domain $D_{s_{c}}$.
A-GEM limits gradient updates to minimize memory loss, improving computational complexity over GEM.
AdapterCL adds domain-specific adapters [18]; the adapter with the lowest action perplexity is selected during evaluation.
Method I-CVLN D-CVLN
[…]
200 23.8 26.3 10.2 50 4.9
PerpR 500 23.8 26.4 9.8 100 6.2
1000 28.7 31.1 8.7 200 6.0
200 25.8 29.2 9.2 50 5.3
ESR 500 27.7 31.1 8.8 100 5.5
1000 29.8 34.4 9.1 200 5.9
Table 3: Comparative analysis of the impact of replay memory size on the performance for rehearsal-based methods. The table shows that larger memory sizes improve performance, indicating a positive correlation between buffer size and agent performance.
5.2 Experiment Setups
For I-CVLN, we use VLN-BERT [17] with Oscar [26] as the backbone, applying CL methods to VLN training. The rehearsal-based approach uses a fixed replay memory of 500, about 3% of the dataset, equally divided among scene domains. AdapterCL employs 20 adapters, matching the scene domains in I-CVLN, and adapters can increase with more domains. Three curricula with randomized scene domain orders are generated using distinct seeds, and results are averaged across them. For non-comparative analyses, results from a single curriculum are reported. For D-CVLN, we use a randomly initialized HAMT [11] as the backbone, applying CL methods to its VLN training. The replay memory size is set to 100, about 3% of the dataset, with 11 adapters fixed for AdapterCL. Three distinct curricula, each covering 11 scene domains, are configured, and results are averaged across these curricula.
5.3 Comparative Results
The results in Table 2 provide a comparative analysis of various methods on I-CVLN and D-CVLN, highlighting the impact of continual learning (CL) strategies in overcoming catastrophic forgetting during sequential training.
Performance Bounds: Vanilla and Joint provide performance bounds, where Vanilla reflects a lower bound due to a lack of mechanisms to mitigate forgetting, and Joint represents an upper bound with full access to all data simultaneously. The large performance gap between these two methods underscores the difficulty of applying continual learning to CVLN, where each new scene domain introduces unique challenges that increase the risk of forgetting prior knowledge.
Performance of Traditional CL Approaches: Standard CL methods such as L2, A-GEM, and AdapterCL fall short of even the Vanilla baseline, indicating that traditional CL strategies struggle with the complex, dynamic environments of I-CVLN and D-CVLN. This result demonstrates that while these methods can help manage forgetting in simpler tasks, they lack the robustness required for the sequential, environment-dependent learning in CVLN.
Effectiveness of Rehearsal-Based Approach: RandR achieves a noticeable improvement over other traditional methods, showcasing the importance of rehearsal for knowledge retention in sequential training. However, its performance still falls short of Joint, suggesting that further refinement in rehearsal strategies could yield more significant gains.
Proposed Methods – PerpR and ESR: Our methods, PerpR and ESR, build upon the strengths of RandR and outperform other CL methods across all metrics. These results indicate that our proposed approaches are effective in retaining previously learned knowledge and adapting to new scene domains. Notably, ESR excels on I-CVLN, while PerpR achieves the best results on D-CVLN, suggesting that the optimal method may vary depending on the specific evaluation setting. These findings highlight the adaptability and robustness of PerpR and ESR in handling the unique demands of CVLN tasks.
Our proposed methods provide a significant improvement in mitigating catastrophic forgetting and adapting to new environments, making them more suitable for complex, multi-domain CVLN tasks than traditional CL methods.
5.4 Analysis
5.4.1 Memory Size Analysis
The proposed methods are rehearsal-based methods that use replay memory. Since the performance of these methods can vary depending on the size of the replay memory, we examined the performance differences of rehearsal-based methods according to the size of their replay memory in Table 3. PerpR and ESR perform better as the replay memory size increases on both datasets. This demonstrates that larger memory can help retain previously learned knowledge by storing a wider variety of data samples.
5.4.2 Stability-Plasticity Trade-off Analysis
In CVLN, the agents need to have memory stability and learning plasticity. Memory stability ensures retention of previously acquired knowledge, while learning plasticity enables the efficient assimilation of new information. These attributes are often in trade-off, known as the stability-plasticity dilemma [33]. Stability (S) is defined as the agents’ average performance on previously encountered scene domains after training on a new one, while Plasticity (P) represents the average initial performance in newly encountered domains. Catastrophic forgetting occurs when plasticity exceeds stability. We evaluate the stability-plasticity trade-off [37], the harmonic mean of S and P, visualized in Figure 5. ESR achieves a better trade-off in I-CVLN than RandR and Vanilla, with greater plasticity and stability, while PerpR performs similarly in D-CVLN.
Ablation I-CVLN D-CVLN
[…]
PerpR $\textbf{26.1}\pm 1.6$ $\textbf{28.9}\pm 1.8$ $\textbf{9.8}\pm 0.2$ $\textbf{6.1}\pm 0.4$
PerpR-Rev. $24.6\pm 1.1$ $27.1\pm 1.2$ $10.2\pm 0.3$ $5.4\pm 0.2$
Table 4: Impact of memory update reversal on performance in PerpR.
Figure 5: Stability-plasticity trade-off comparison in I-CVLN and D-CVLN: We calculate stability and plasticity for agents after learning 10 scene domains.
5.5 Ablation Study
5.5.1 Reversing PerpR memory update
Table 4 shows the results of reversing the replay memory update in PerpR. PerpR is developed based on the assumption that episodes with higher action perplexity would be more beneficial for future learning. The results confirm the validity of PerpR’s underlying assumption. PerpR’s performance degrades when replay memory updates are reversed by storing low action perplexity episodes and deleting high ones.
5.5.2 ESR loss ablation
Ablation I-CVLN D-CVLN
[…]
$\mathcal{L}_{ESR\_total}$ 27.7 31.1 8.8 5.5
- $\mathcal{L}_{M}$ 26.0 27.9 9.7 5.4
- $\mathcal{L}_{ESR}$ 22.7 24.9 10.4 5.1
Table 5: Ablation study on ESR for I-CVLN and D-CVLN. This table shows the effects of removing specific loss components on the overall performance.
Table 5 shows the results of an ablation study on the ESR loss in Equation 7. The removal of each of the two losses shows decreased performance across all datasets. Notably, the removal of $\mathcal{L}_{ESR}$ resulted in a greater performance drop than the removal of $\mathcal{L}_{M}$. This confirms that leveraging the action distribution predictions from previous models helps mitigate the catastrophic forgetting problem in CVLN agents. Additionally, we observe a synergistic effect when using both types of losses obtained through replay memory.
6 Conclusions and Future Work
In this paper, we introduced the Continual Vision-and-Language Navigation (CVLN) paradigm to enable VLN agents to learn in changing environments without forgetting previous knowledge. We proposed two settings, I-CVLN and D-CVLN, to account for various forms of natural language instruction. Additionally, we present two simple yet effective methods—Perplexity Replay and Episodic Self-Replay—that use rehearsal mechanisms to mitigate catastrophic forgetting and are applicable to tasks like CVLN, which involve sequential decision-making. Our experiments showed that these methods outperform existing continual learning approaches, highlighting the importance of effective replay memory in CVLN. For future work, we plan to extend CVLN to continuous environments and design new settings and datasets specifically for VLN continual learning, beyond existing dataset-based benchmarks.

### v3

@@
-Continual Vision-and-Language Navigation
+
 Abstract
-In developing Vision-and-Language Navigation (VLN) agents that navigate to a destination using natural language instructions and visual cues, current studies largely assume a train-once-deploy-once strategy.
-We argue that this kind of strategy is less realistic, as deployed VLN agents are expected to encounter novel environments continuously through their lifetime.
-To facilitate more realistic setting for VLN agents, we propose Continual Vision-and-Language Navigation (CVLN) paradigm for agents to continually learn and adapt to changing environments.
-In CVLN, the agents are trained and evaluated incrementally across multiple scene domains (i.e., environments).
-We present two CVLN learning setups to consider diverse forms of natural language instructions: Initial-instruction based CVLN, focused on navigation via initial-instruction interpretation, and dialogue-based CVLN, designed for navigation through dialogue with other agents.
-We introduce two simple yet effective baseline methods, tailored to the sequential decision-making needs of CVLN: Perplexity Replay (PerpR) and Episodic Self-Replay (ESR), both employing a rehearsal mechanism.
-PerpR selects replay episodes based on episode difficulty, while ESR stores and revisits action logits from individual episode steps during training to refine learning.
-Experimental results indicate that while existing continual learning methods are insufficient for CVLN, PerpR and ESR outperform the comparison methods by effectively utilizing replay memory.
+Developing Vision-and-Language Navigation (VLN) agents typically assumes a train-once-deploy-once strategy, which is unrealistic as deployed agents continually encounter novel environments.
+To address this, we propose the Continual Vision-and-Language Navigation (CVLN) paradigm, where agents learn and adapt incrementally across multiple scene domains.
+CVLN includes two setups: Initial-instruction based CVLN for instruction-following, and Dialogue-based CVLN for dialogue-guided navigation.
+We also introduce two simple yet effective baselines for sequential decision-making: Perplexity Replay (PerpR), which replays difficult episodes, and Episodic Self-Replay (ESR), which stores and revisits action logits during training.
+Experiments show that existing continual learning methods fall short for CVLN, while PerpR and ESR achieve better performance by efficiently utilizing replay memory.
 1 Introduction
 One of the desired attributes of robotic agents is their ability to learn and adapt to new environments without losing knowledge of previous ones.
-In practice, robots deployed in the human world are required to perform a variety of tasks in complex and changing environments.
+In practice, robots deployed in the real world continually encounter unseen environments and are required to perform a variety of tasks in them.
 Vision-and-Language Navigation (VLN) focuses on agents that navigate using natural language instructions and visual cues, promising for applications like robotic assistants.
 Despite recent advancements, most VLN research follows a train-once-deploy-once strategy.
-In real-world applications, environments are not static; they evolve over time with new buildings, renovations, or changes in layout.
 A deployed VLN agent must promptly adapt whenever the environment changes throughout its lifetime.
-Finetuning on new environments might seem practical, but it can cause the agent’s performance in previously learned contexts to deteriorate, a challenge known as catastrophic forgetting [30].
-To address this challenge, the agent needs the capability for Continual Learning (CL) [42], which enables it to learn and adapt to new circumstances while retaining its existing knowledge.
+Finetuning on new environments might seem practical, but it can cause the agent’s performance in previously learned contexts to deteriorate, a challenge known as catastrophic forgetting [McCloskey and Cohen(1989)].
+To address this challenge, the agent needs the capability for Continual Learning (CL) [Thrun and Pratt(2012)], which enables it to learn and adapt to new circumstances while retaining its existing knowledge.
 In this paper, we propose Continual Vision-and-Language Navigation (CVLN) paradigm, which trains and evaluates VLN agents with consideration of the CL ability.
@@
 Figure 1 illustrates an overview of CVLN.
-In CVLN, the agent incrementally learns multiple scene domains, and as shown in Figure 1, there are substantial differences in visual distributions among scene domains.
-These distributional discrepancies can lead to catastrophic forgetting when the domains are learned incrementally.
 The CVLN supports two settings to handle different types of natural language instructions: Initial-instruction based CVLN (I-CVLN) and Dialogue based CVLN (D-CVLN).
@@
 Furthermore, we conducted an ablation study, analyzed the effects of varying replay memory sizes, and examined the stability-plasticity trade-off.
-Figure 1: Frechet Inception Distance[16] heatmaps illustrating the distribution shifts between visual features [47] across 10 scene domains for the I-CVLN and D-CVLN.
-To summarize our contributions:
-  • We propose a Continual Vision-and-Language Navigation (CVLN) paradigm to develop VLN agents capable of continual learning, introducing two settings— I-CVLN (instruction-based) and D-CVLN (dialogue-based)—for comprehensive evaluation across diverse natural language instruction types.
-  • We propose two simple yet effective CVLN methods, PerpR and ESR, which enhance continual learning through rehearsal mechanisms, applicable to tasks like CVLN that involve sequential decision-making.
-  • In experiments, PerpR and ESR demonstrate their effectiveness in the CVLN paradigm, achieving improved performance over comparative methods in both I-CVLN and D-CVLN settings.
-Figure 2: Comparison between (a) VLN and (b) CVLN.
-In VLN, the agent is trained within fixed scene domains and then evaluated on unseen scene domains.
-In contrast, the agent in CVLN is trained within a new scene domain sequentially and and evaluated on those scene domains along with previously encountered ones.
+Figure 1: Comparison between (a) VLN and (b) CVLN: VLN trains agents on fixed environments and evaluates them on unseen ones, while CVLN trains agents sequentially on new environments and evaluates them on both previously encountered and newly learned environments.
 2 Related Work
@@
 Continual Learning (CL) focuses on training models to acquire new knowledge over time while retaining previously learned information.
-CL methods have generally been extensively studied for simple tasks [13] in computer vision.
+CL methods have generally been extensively studied for simple tasks [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei] in computer vision.
 This field includes several CL scenarios.
-Instance-Incremental Learning trains on batches of samples for the same task [27].
-Domain-Incremental Learning (DIL) deals with tasks having the same labels but different input distributions, with no domain identity during inference [19].
-Task-Incremental and Class-Incremental Learning involve different label spaces, with varying use of domain identity [19].
-Task-Free Continual Learning handles separate label spaces without domain identity [1], and Online Continual Learning processes sequential data streams in real-time [2].
+Instance-Incremental Learning trains on batches of samples for the same task [Lomonaco and Maltoni(2017)].
+Domain-Incremental Learning (DIL) deals with tasks having the same labels but different input distributions, with no domain identity during inference [Hsu et al.(2018)Hsu, Liu, Ramasamy, and Kira].
+Task-Incremental and Class-Incremental Learning involve different label spaces, with varying use of domain identity [Hsu et al.(2018)Hsu, Liu, Ramasamy, and Kira].
+Task-Free Continual Learning handles separate label spaces without domain identity [Aljundi et al.(2019a)Aljundi, Kelchtermans, and Tuytelaars], and Online Continual Learning processes sequential data streams in real-time [Aljundi et al.(2019b)Aljundi, Lin, Goujaud, and Bengio].
 CVLN fits within DIL, providing scene domain identities during training but not during testing, requiring generalization without explicit domain knowledge at evaluation.
-There are three main approaches for CL: Regularization [22], which adds terms to prevent the loss of prior knowledge; Rehearsal [34, 9, 8, 7], which stores and reuses data from past tasks; and Architectural [36, 29], which introduces task-specific parameters to the agent’s architecture.
+There are three main approaches for CL: Regularization [Kirkpatrick et al.(2017)Kirkpatrick, Pascanu, Rabinowitz, Veness, Desjardins, Rusu, Milan, Quan, Ramalho, Grabska-Barwinska, et al.], which adds terms to prevent the loss of prior knowledge; Rehearsal [Rebuffi et al.(2017)Rebuffi, Kolesnikov, Sperl, and Lampert, Chaudhry et al.(2018)Chaudhry, Ranzato, Rohrbach, and Elhoseiny, Buzzega et al.(2020)Buzzega, Boschini, Porrello, Abati, and Calderara, Boschini et al.(2022)Boschini, Bonicelli, Buzzega, Porrello, and Calderara], which stores and reuses data from past tasks; and Architectural [Rusu et al.(2016)Rusu, Rabinowitz, Desjardins, Soyer, Kirkpatrick, Kavukcuoglu, Pascanu, and Hadsell, Madotto et al.(2020)Madotto, Lin, Zhou, Moon, Crook, Liu, Yu, Cho, and Wang], which introduces task-specific parameters to the agent’s architecture.
 This paper proposes two Rehearsal-based methods, PerpR and ESR, and evaluates their effectiveness through comparative experiments.
@@
 In VLN, agents navigate realistic 3D indoor environments by following natural language instructions.
-Various datasets like R2R [6], R4R [20], and RxR [24] have been developed, focusing on navigation based on initial instructions that fully describe the path.
-Other tasks, such as CVDN [41] and HANNA [32], involve agents obtaining navigation information through dialogue with an oracle or other agents.
-Additionally, datasets like Touchdown [10] and StreetLearn [31] have expanded VLN research to outdoor environments.
-IVLN [23] present a paradigm where agents maintain memory across episodes to navigate a sequence of episodes.
-In the field of VLN, research focuses on improving agent architecture [17, 11, 14, 4], enhancing visual-text alignment through auxiliary losses [28, 46, 48], using data augmentation [40, 45, 25, 21], and extensive pre-training [15, 11, 12, 3] to enhance generalization.
-In this paper, we adopted VLN-BERT [17] and HAMT [11] as the backbone architectures.
+Various datasets like R2R [Anderson et al.(2018b)Anderson, Wu, Teney, Bruce, Johnson, Sünderhauf, Reid, Gould, and Van Den Hengel], R4R [Jain et al.(2019)Jain, Magalhaes, Ku, Vaswani, Ie, and Baldridge], and RxR [Ku et al.(2020)Ku, Anderson, Patel, Ie, and Baldridge] have been developed, focusing on navigation based on initial instructions that fully describe the path.
+Other tasks, such as CVDN [Thomason et al.(2020)Thomason, Murray, Cakmak, and Zettlemoyer] and HANNA [Nguyen and Daumé III(2019)], involve agents obtaining navigation information through dialogue with an oracle or other agents.
+Additionally, datasets like Touchdown [Chen et al.(2019)Chen, Suhr, Misra, Snavely, and Artzi] and StreetLearn [Mirowski et al.(2019)Mirowski, Banki-Horvath, Anderson, Teplyashin, Hermann, Malinowski, Grimes, Simonyan, Kavukcuoglu, Zisserman, et al.] have expanded VLN research to outdoor environments.
+IVLN [Krantz et al.(2023)Krantz, Banerjee, Zhu, Corso, Anderson, Lee, and Thomason] present a paradigm where agents maintain memory across episodes to navigate a sequence of episodes.
+In the field of VLN, research focuses on improving agent architecture [Hong et al.(2020)Hong, Wu, Qi, Rodriguez-Opazo, and Gould, Chen et al.(2021)Chen, Guhur, Schmid, and Laptev, Fried et al.(2018)Fried, Hu, Cirik, Rohrbach, Andreas, Morency, Berg-Kirkpatrick, Saenko, Klein, and Darrell, An et al.(2024)An, Wang, Wang, Wang, Huang, He, and Wang], enhancing visual-text alignment through auxiliary losses [Ma et al.(2019)Ma, Lu, Wu, AlRegib, Kira, Socher, and Xiong, Wang et al.(2019)Wang, Huang, Celikyilmaz, Gao, Shen, Wang, Wang, and Zhang, Zhu et al.(2020)Zhu, Zhu, Chang, and Liang], using data augmentation [Tan et al.(2019)Tan, Yu, and Bansal, Wang et al.(2022)Wang, Montgomery, Orbay, Birodkar, Faust, Gur, Jaques, Waters, Baldridge, and Anderson, Li et al.(2022)Li, Tan, and Bansal, Kamath et al.(2023)Kamath, Anderson, Wang, Koh, Ku, Waters, Yang, Baldridge, and Parekh], and extensive pre-training [Hao et al.(2020)Hao, Li, Li, Carin, and Gao, Chen et al.(2021)Chen, Guhur, Schmid, and Laptev, Chen et al.(2022)Chen, Guhur, Tapaswi, Schmid, and Laptev, An et al.(2023)An, Qi, Li, Huang, Wang, Tan, and Shao] to enhance generalization.
+In this paper, we adopted VLN-BERT [Hong et al.(2020)Hong, Wu, Qi, Rodriguez-Opazo, and Gould] and HAMT [Chen et al.(2021)Chen, Guhur, Schmid, and Laptev] as the backbone architectures.
 Existing VLN datasets and methods lack considerations for CL, highlighting the need for scalable strategies that enable agents to learn efficiently to diverse, changing environments without costly retraining models from scratch.
 3 Continual Vision-and-Lauguage Navigation
-3.1 Formulation of VLN
-The existing VLN agents are trained on fixed training data and evaluated on episodes from scenes that were not encountered during training.
-The objective of VLN agents is to minimize the following loss:
-$\mathcal{L}_{D}\triangleq\mathbb{E}_{(I,A^{*})\sim D}[\ell(\pi_{\theta}(V,I),A^{*})]$ (1)
-where $\theta$ specifies the parameters for the VLN agent, $\pi$.
-The dataset $D$ is composed of episodes, which include navigation instructions $I$ and their corresponding ground-truth trajectory $A^{*}=\{a_{1}^{*},a_{2}^{*},\ldots,a_{N}^{*}\}$.
-The set $V=\{v_{1},v_{2},\ldots,v_{N}\}$ represents the sequence of visual observations from the environment, each associated with a step determined by a predicted action, with $N$ denoting the total number of steps per episode.
-The agent $\pi_{\theta}(V,I)$ determines the actions based on $V$ and $I$ sequentially.
-The navigation imitation loss function $\ell$ then compares this behavior to the ground-truth trajectory $A^{*}$ to evaluate the discrepancy between them.
-We use cross entropy loss for $\ell$.
-3.2 Formulation of CVLN
-A scene domain includes episodes in multiple indoor environments, and each episode comprises a natural language navigation instruction and its corresponding trajectory.
-CVLN agent learns scene domains incrementally.
-The agent is evaluated on the incrementally learned scene domains.
-The evaluation episodes belong to the learned scene domains but are episodes that have not been seen during training.
-During this process, the agent must learn the current scene domain without forgetting the knowledge about previously learned scene domains.
-This learning approach is more suitable for real-world agents that encounter and need to learn new scene domains continually.
-Formally, a CVLN is divided into $S$ scene domains; during each scene domain $s\in\{1,\ldots,S\}$, an episode $\epsilon$, navigation instruction $I$ and their corresponding ground-truth trajectory $A^{*}$ are drawn from an independent and identically distributed (i.i.d.) distribution $D_{s}$.
-The objective of agents in CVLN is as follows:
+3.1 Formulation of VLN and CVLN
+In VLN, agents are trained on fixed data and evaluated on episodes from unseen scenes.
+Given a dataset $D$ of episodes with navigation instructions $I$ and ground-truth trajectories $A^{*}=\{a_{1}^{*},\ldots,a_{N}^{*}\}$, the agent $\pi_{\theta}$ sequentially predicts actions based on visual observations $V=\{v_{1},\ldots,v_{N}\}$ and instructions $I$.
+The objective minimizes the navigation imitation loss $\ell$, typically cross-entropy:
 […]
-$\quad\text{where}\ \mathcal{L}_{D_{s}}\triangleq\mathbb{E}_{(I,A^{*})\sim D_{s}}[\ell(\pi_{\theta}(V,I),A^{*})]$
+CVLN extends this by introducing continual learning across $S$ scene domains.
+In each domain $s$, episodes are drawn from distribution $D_{s}$, and the agent must learn new domains without forgetting previous ones.
+The learning objective becomes:
+[…]
+This continual learning setting better reflects real-world scenarios where agents encounter and must adapt to new environments over time.
 Figure 2: Comparison between I-CVLN and D-CVLN.
@@
 where an agent $\pi$, with parameters $\theta$, is optimized on one scene domain at a time in a sequential manner. $s_{c}$ represents the current scene domain.
-3.3 Initial-Instruction based CVLN
-In I-CVLN, agents receive initial instructions outlining the entire navigation path.
-As the R2R [6] and RxR [24] datasets provide initial instruction-based navigation episodes, we construct our continual learning benchmark for I-CVLN on top of these datasets, using only english epidodes.
-The training split from the original datasets is used for I-CVLN training, while the val-seen split, which contain episodes from the same scenes but not seen during training, is used for I-CVLN validation.
-We grouped episodes from the same scenes across both datasets into 20 scene domains, each containing over 600 episodes for training.
-Datasets Split Episode Scene Domain
-I-CVLN Train 15700 20
-Valid 1563 20
-D-CVLN Train 3737 11
-Valid 382 11
-Table 1: Dataset statistics for I-CVLN and D-CVLN. The table outlines the number of episodes and scene domains available in the training (Train) and validation (Valid) splits for each dataset.
-3.4 Dialogue based CVLN
-In D-CVLN, agents navigate from dialogue history.
-As the CVDN [41] dataset includes dialogue data between a follower and an oracle paired with corresponding trajectories, we build benchmark for D-CVLN on top of it.
-Due to the limited episodes per scene in CVDN, we grouped four scenes into each scene domain, resulting in 11 domains for training.
-These scene domains are learned sequentially.
-The training split of CVDN is used for D-CVLN training, while the val-seen split, containing evaluation episodes for each domain but not seen during training, is used for D-CVLN validation.
-Figure 3 illustrates a comparison between I-CVLN and D-CVLN.
-3.5 Evaluation Protocol for CVLN
-The evaluation of CVLN agents focuses on assessing their ability to retain previously learned knowledge while effectively adapting to new scene domain.
-For CVLN evaluation, the Average Metric is used to evaluate agent performance across different scene domains.
-The average metric is defined as follows:
+3.2 I-CVLN and D-CVLN Benchmarks
+We design two CVLN benchmarks: Initial-Instruction based CVLN (I-CVLN) and Dialogue-based CVLN (D-CVLN), constructed from existing VLN datasets.
+In I-CVLN, agents receive an initial instruction describing the full navigation path.
+Using the R2R [Anderson et al.(2018b)Anderson, Wu, Teney, Bruce, Johnson, Sünderhauf, Reid, Gould, and Van Den Hengel] and RxR [Ku et al.(2020)Ku, Anderson, Patel, Ie, and Baldridge] datasets, we construct the benchmark with English-only episodes.
+The training split from these datasets provides 15,700 episodes, while validation uses 1,563 unseen episodes from the same scenes (val-seen split).
+Episodes are grouped into 20 scene domains, each containing over 600 training episodes.
+In D-CVLN, agents navigate based on dialogue history.
+This benchmark is built using the CVDN [Thomason et al.(2020)Thomason, Murray, Cakmak, and Zettlemoyer] dataset, which includes dialogues between a follower and an oracle paired with corresponding trajectories.
+Due to limited episodes per scene, four scenes are grouped into each scene domain, resulting in 11 domains.
+The training split includes 3,737 episodes, and validation uses 382 unseen episodes from the same domains.
+In both benchmarks, scene domains are learned sequentially, aligning with the continual learning setting.
+3.3 Evaluation of CVLN
+CVLN evaluation measures an agent’s ability to retain prior knowledge while adapting to new scene domains. We use the Average Metric, defined as:
 $\frac{1}{S}\sum_{i=1}^{S}R_{S,i}$ (3)
-where $R_{S,i}$ indicates the agent’s performance on the $i$th domain after training the $S$th scene domain.
-The metric evaluates the agent’s knowledge retention across prior scene domains and its adaptability to new ones.
-For I-CVLN, we use standard VLN metrics [5] $R_{S,i}$ in Equation 3: Success Rate (SR), Success weighted by inverse Path Length (SPL), and Navigation Error (NE).
-For D-CVLN, we use the main metric from CVDN [41], $R_{S,i}$ in Equation 3, which is Goal Progress (GP) in meters.
+where $R_{S,i}$ is the agent’s performance on the $i$th domain after learning the $S$th domain.
+For I-CVLN, $R_{S,i}$ uses standard VLN metrics [Anderson et al.(2018a)Anderson, Chang, Chaplot, Dosovitskiy, Gupta, Koltun, Kosecka, Malik, Mottaghi, Savva, et al.]: Success Rate (SR), Success weighted by inverse Path Length (SPL), and Navigation Error (NE).
+For D-CVLN, we use Goal Progress (GP) in meters from CVDN [Thomason et al.(2020)Thomason, Murray, Cakmak, and Zettlemoyer].
 4 Methods
 Rehearsal-based CL approaches mitigate catastrophic forgetting by using a fixed-size replay memory $M$ to store and revisit episodes from previous domains.
-After training each task, $M$ is updated by deleting an equal number of old episodes and adding new ones from the latest domain.
+After training each domain, $M$ is updated by deleting an equal number of old episodes and adding new ones from the latest domain.
 Effective replay memory construction and usage are crucial for maintaining performance across domains.
@@
 Figure 3 provides an overview of PerpR and ESR.
-Figure 4: Overview of Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) for CVLN agents.
-(a) PerpR prioritizes challenging episodes in the agent’s memory, optimizing for high Action Perplexity (AP).
-(b) ESR enables the agent to self-replay using past optimal behaviors.
+Figure 3: Overview of Perplexity Replay (PerpR) and Episodic Self-Replay (ESR) for CVLN: (a) PerpR prioritizes challenging episodes with high Action Perplexity (AP), and (b) ESR enables self-replay using past behaviors.
 4.1 Perplexity Replay
-To construct replay memory for continual learning, random sampling [44, 35] is often employed.
-However, this approach fails to consider the importance of samples in recovering past knowledge, resulting in suboptimal replay memory composition.
-Existing methods [39, 2] have attempted to address this by utilizing gradient-based or optimizable approaches to construct more effective replay memory.
-Nevertheless, these methods are typically tailored to simple tasks such as image classification and are not suitable for tasks that require sequential action prediction, such as CVLN.
-Algorithm 1 Perplexity Replay Memory Update
-1:  Input: Replay memory $M$, replay memory size $\lvert M\rvert$, set of episodes from the last scene domain $E_{\text{last}}$, agent parameters $\theta$, the number of trained scene domains $s$
-2:  Output: Updated replay memory $M$
-3:  Calculate Action Perplexity (AP) for each episode in $E_{\text{last}}$ and $M$
-[…]
-[…]
-6:  $n_{s}\leftarrow\lvert M\rvert/(s+1)$
-7:  for $s=1$ to $s$ do
-8:     Remove episodes with the lowest AP from $M$ for domain $s$ until only $n_{s}$ episodes remain
-9:  end for
-10:  Add episodes from $E_{\text{last}}$ to $M$ until $\lvert M\rvert$ equals fixed size
-11:  return  $M$
-To overcome these limitations, we propose Perplexity Replay (PerpR), which constructs the replay memory by quantifying the difficulty of VLN episodes.
-A key innovation of PerpR is that the model self-evaluates the difficulty of each episode using Action Perplexity (AP), without relying on external metrics or annotations.
-By self-assessing its own uncertainty through AP, the model identifies challenging episodes where it is less confident.
-PerpR then prioritizes these difficult episodes in the replay memory, enabling the agent to focus on experiences that require further learning and thus enhancing the overall learning process.
+Random sampling [Vitter(1985), Riemer et al.(2018)Riemer, Cases, Ajemian, Liu, Rish, Tu, and Tesauro] is often used to construct replay memory for continual learning but overlooks sample importance for recovering past knowledge, leading to suboptimal memory composition.
+Existing methods [Shim et al.(2021)Shim, Mai, Jeong, Sanner, Kim, and Jang, Aljundi et al.(2019b)Aljundi, Lin, Goujaud, and Bengio] address this using gradient-based or optimizable approaches, but they are mainly designed for simple tasks like image classification and are unsuitable for sequential action prediction required in CVLN.
+To address these limitations, we propose Perplexity Replay (PerpR), which builds replay memory by quantifying episode difficulty.
+PerpR introduces Action Perplexity (AP) to let the model self-evaluate episode difficulty without external metrics or annotations.
+Using AP, the model identifies challenging episodes with high uncertainty and prioritizes them in replay memory, focusing learning on harder experiences and improving overall performance.
 Assuming episodes with high AP indicate insufficient agent learning, PerpR ensures that the replay memory is populated with experiences most beneficial for the agent’s continued improvement.
@@
 $\mathcal{L}_{PerpR}=\mathcal{L}_{D_{s_{c}}}+\mathcal{L}_{M}$ (5)
-$\mathcal{L}_{D_{s_{c}}}$ $\triangleq\mathbb{E}_{(I,A^{*})\sim D_{s_{c}}}[\ell(\pi_{\theta}(V,I),A^{*})]$
-$\mathcal{L}_{M}$ $\triangleq\mathbb{E}_{(I,A^{*})\sim M}[\ell(\pi_{\theta}(V,I),A^{*})]$
+[…]
+[…]
 where $\mathcal{L}_{D_{s_{c}}}+\mathcal{L}_{M}$ represents the sum of the losses over the current scene domain $D_{s_{c}}$ and the replay memory $M$.
-The memory update procedure for PerpR is described in Algorithm 1.
+The memory update procedure for PerpR is provided in the supplementary material (Section 4.1).
 4.2 Episodic Self-Replay
-To effectively utilize the samples in the replay memory, prior works [34, 8, 7, 43] have employed logit distillation methods, which were designed for continual learning in non-sequential tasks such as image classification or object detection.
-However, they are not directly applicable to CVLN due to its sequential action prediction in dynamic environments.
-In CVLN, agents must perform sequences of actions where observations depend on previous actions, involving temporal state changes and decision-making processes.
-Existing logit distillation methods focus on single-step predictions, limiting their effectiveness in learning or generalizing complex sequential action patterns.
+To leverage replay memory effectively, prior works [Rebuffi et al.(2017)Rebuffi, Kolesnikov, Sperl, and Lampert, Buzzega et al.(2020)Buzzega, Boschini, Porrello, Abati, and Calderara, Boschini et al.(2022)Boschini, Bonicelli, Buzzega, Porrello, and Calderara, Verwimp et al.(2022)Verwimp, Yang, Parisot, Lanqing, McDonagh, Pérez-Pellitero, De Lange, and Tuytelaars] used logit distillation for non-sequential tasks like image classification.
+However, these methods are unsuitable for CVLN, which requires sequential action prediction.
+CVLN agents must handle temporally dependent observations and decision-making, while existing distillation methods focus only on single-step predictions, limiting their ability to learn complex action sequences.
 To address this limitation, we extended these approaches for application to CVLN episodes.
@@
 When learning from episodes in the replay memory, the CVLN agent utilizes these stored logits to train itself to follow its past action distribution at each step.
-To reinforce consistency, we applied student forcing [38] during training by using the stored logits as the next action for subsequent steps, rather than the predicted logits.
+To reinforce consistency, we applied student forcing [Sauder et al.(2020)Sauder, Hu, Che, Mordido, Yang, and Meinel] during training by using the stored logits as the next action for subsequent steps, rather than the predicted logits.
 This encourages adherence to prior action patterns and aligns the learning process with the sequential nature of CVLN.
-Algorithm 2 Episodic Self-Replay
-0:  Dataset $D_{s}$, Memory $M$, Learning rate $\eta$, Agent $f_{\theta}$, $\text{GetObservation}(t)$, $\text{ActionExecuter}(a_{t})$, $\lambda_{1}$, $\lambda_{2}$
-1:  Sample $(I,A^{*})\sim D_{s}$
-2:  $\mathcal{L}_{D_{c}}\leftarrow 0$, $Z\leftarrow[]$, $t\leftarrow 1$
-3:  $v_{t}\leftarrow\text{GetObservation}(t)$;
-4:  while $a_{t}$ does not predict stop do
-5:     $z_{t}\leftarrow f_{\theta}(I,v_{t})$; $Z.\text{append}(z_{t})$
-6:     $a_{t}\leftarrow\arg\max(z_{t})$
-7:     $a_{t}^{*}\leftarrow A^{*}(t)$ //Get true action
-8:     $\mathcal{L}_{D_{c}}\leftarrow\mathcal{L}_{D_{c}}+\ell(a_{t},a_{t}^{*})$
-9:     $v_{t+1}\leftarrow\text{ActionExecuter}(a_{t})$
-10:     $t\leftarrow t+1$
-11:  end while
-12:  $\mathcal{L}_{D_{c}}\leftarrow\mathcal{L}_{D_{c}}/t$
-13:  Update $M\leftarrow\text{Random}(M,(I,A^{*},Z))$
-14:  Sample $(I^{\prime},A^{\prime*},Z^{\prime*})\sim M$
-15:  $\mathcal{L}_{ESR}\leftarrow 0$, $\mathcal{L}_{M}\leftarrow 0$, $t\leftarrow 1$
-16:  $v_{t}\leftarrow\text{GetObservation}(t)$;
-17:  while $a_{t}$ does not predict stop do
-18:     $z_{t}\leftarrow f_{\theta}(I^{\prime},v_{t})$
-19:     $a_{t}\leftarrow\arg\max(z_{t})$
-20:     $a_{t}^{\prime*}\leftarrow A^{\prime*}(t)$ //Get true action
-21:     $\mathcal{L}_{M}\leftarrow\mathcal{L}_{M}+\ell(a_{t},a_{t}^{\prime*})$
-22:     $z_{t}^{\prime*}\leftarrow Z^{\prime*}(t)$ //Get replay logit
-23:     $\mathcal{L}_{ESR}\leftarrow\mathcal{L}_{ESR}+\|z_{t}-z_{t}^{\prime*}\|^{2}_{F}$//Logit distillation
-24:     $a_{t}^{*}\leftarrow\arg\max(z_{t}^{\prime*})$ //Get replay logit’s action
-25:     $v_{t+1}\leftarrow\text{ActionExecuter}(a_{t}^{*})$
-26:     $t\leftarrow t+1$
-27:  end while
-28:  $\mathcal{L}_{ESR}\leftarrow\mathcal{L}_{ESR}/t$
-29:  $\mathcal{L}_{M}\leftarrow\mathcal{L}_{M}/t$
-30:  $\mathcal{L}_{ESR\_{total}}\leftarrow\mathcal{L}_{D_{c}}+\lambda_{1}\mathcal{L}_{ESR}+\lambda_{2}\mathcal{L}_{M}$
-31:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{ESR\_{total}}$
-We define the agent parameters immediately after learning the scene domain $s$ as $\theta_{s}^{*}$.
-A function $f$ outputs the logits for action candidates at each sequential step of navigation.
-Unlike other rehearsal-based methods, ESR stores the sequential action logits $Z\triangleq\{z_{n}\}_{n=1}^{N}\triangleq f_{\theta_{s}^{*}}(V,I)$, corresponding to each step in the sequence, when updating the replay memory $M$.
-ESR’s replay memory uses random sampling [44] to select episodes to add and delete.
+[…]
 The ESR loss function can be formulated as follows:
-$\mathcal{L}_{ESR}\triangleq\mathbb{E}_{(I,Z)\sim M}\left[\sum_{n=1}^{N}\|z_{n}-f_{\theta}(V_{n},I_{n})\|^{2}_{F}\right]$ (6)
-Method I-CVLN D-CVLN
 […]
-1. Vanilla $23.1\pm 0.8$ $25.6\pm 0.6$ $10.5\pm 0.5$ $5.5\pm 0.4$
-2. Joint $40.1\pm 0.2$ $42.6\pm 0.1$ $7.1\pm 0.3$ $8.1\pm 0.1$
-3. L2 $13.3\pm 0.2$ $14.6\pm 0.2$ $12.7\pm 0.3$ $4.3\pm 0.4$
-4. AdapterCL $7.8\pm 0.3$ $9.1\pm 0.6$ $13.9\pm 0.8$ $4.3\pm 1.2$
-5. AGEM $4.6\pm 2.8$ $4.5\pm 2.2$ $13.0\pm 0.2$ $2.7\pm 0.9$
-6. RandR $24.9\pm 0.6$ $28.0\pm 0.7$ $10.0\pm 0.5$ $5.7\pm 0.3$
-7. PerpR (Ours) $\underline{26.1}\pm 1.6$ $\underline{28.9}\pm 1.8$ $\underline{9.8}\pm 0.2$ $\textbf{6.1}\pm 0.4$
-8. ESR (Ours) $\textbf{28.2}\pm 0.4$ $\textbf{31.9}\pm 0.9$ $\textbf{9.2}\pm 0.3$ $\underline{5.8}\pm 0.5$
-Table 2: Comparative results of various methods on I-CVLN and D-CVLN. This table reports the mean and standard deviation values across three distinct learning curricula.
 ESR uses $\mathcal{L}_{ESR}$ to allow agents to replay their behavior distribution during the learning process, which allows them to learn about new scene domains while effectively retaining previously learned knowledge.
@@
 Both $\lambda_{1}$ and $\lambda_{2}$ are set to 0.2.
-Please refer to Algorithm 2 for more details.
+Please refer to supplementary material (Section 4.2) for more details.
 5 Experiments
@@
 L2 Regularization reduces the difference between the current parameter $\theta_{s}$ and the previous $\theta_{s-1}$ by adding the regularization term $\lambda\lVert\theta_{s}-\theta_{s-1}^{*}\rVert^{2}$ to the loss function.
-Random Replay (RandR) uses Reservoir sampling [44] to manage replay memory, integrating selected episodes with current training scene domain $D_{s_{c}}$.
-A-GEM limits gradient updates to minimize memory loss, improving computational complexity over GEM.
-AdapterCL adds domain-specific adapters [18]; the adapter with the lowest action perplexity is selected during evaluation.
-Method I-CVLN D-CVLN
+Random Replay (RandR) uses Reservoir sampling [Vitter(1985)] to manage replay memory, integrating selected episodes with current training scene domain $D_{s_{c}}$.
+A-GEM [Chaudhry et al.(2018)Chaudhry, Ranzato, Rohrbach, and Elhoseiny] limits gradient updates to minimize memory loss, improving computational complexity over GEM [Lopez-Paz and Ranzato(2017)].
+AdapterCL adds domain-specific adapters [Houlsby et al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone, De Laroussilhe, Gesmundo, Attariyan, and Gelly]; the adapter with the lowest action perplexity is selected during evaluation.
+5.2 Experiment Setups
+For I-CVLN, we use VLN-BERT [Hong et al.(2020)Hong, Wu, Qi, Rodriguez-Opazo, and Gould] with Oscar [Li et al.(2020)Li, Yin, Li, Zhang, Hu, Zhang, Wang, Hu, Dong, Wei, et al.] as the backbone and apply CL methods during training.
+The rehearsal-based method uses a fixed replay memory of 500 samples (about 3% of the dataset), equally divided across 20 scene domains.
+AdapterCL employs 20 adapters matching these domains, expandable for more domains.
+Three curricula with different scene orders are generated using distinct seeds, and results are averaged; for non-comparative analyses, a single curriculum is used.
+For D-CVLN, we adopt a randomly initialized HAMT [Chen et al.(2021)Chen, Guhur, Schmid, and Laptev] backbone with CL methods.
+The replay memory is set to 100 samples (about 3% of the dataset), and AdapterCL uses 11 fixed adapters.
+Three curricula covering 11 scene domains are configured, and results are averaged.
 […]
-200 23.8 26.3 10.2 50 4.9
-PerpR 500 23.8 26.4 9.8 100 6.2
-1000 28.7 31.1 8.7 200 6.0
-200 25.8 29.2 9.2 50 5.3
-ESR 500 27.7 31.1 8.8 100 5.5
-1000 29.8 34.4 9.1 200 5.9
-Table 3: Comparative analysis of the impact of replay memory size on the performance for rehearsal-based methods.
-The table shows that larger memory sizes improve performance, indicating a positive correlation between buffer size and agent performance.
-5.2 Experiment Setups
-For I-CVLN, we use VLN-BERT [17] with Oscar [26] as the backbone, applying CL methods to VLN training.
-The rehearsal-based approach uses a fixed replay memory of 500, about 3% of the dataset, equally divided among scene domains.
-AdapterCL employs 20 adapters, matching the scene domains in I-CVLN, and adapters can increase with more domains.
-Three curricula with randomized scene domain orders are generated using distinct seeds, and results are averaged across them.
-For non-comparative analyses, results from a single curriculum are reported.
-For D-CVLN, we use a randomly initialized HAMT [11] as the backbone, applying CL methods to its VLN training.
-The replay memory size is set to 100, about 3% of the dataset, with 11 adapters fixed for AdapterCL.
-Three distinct curricula, each covering 11 scene domains, are configured, and results are averaged across these curricula.
+[…]
+[…]
+2. Joint $40.1\pm 0.2$ $42.6\pm 0.1$ $7.1\pm 0.3$ $8.1\pm 0.1$
+[…]
+4. AdapterCL $7.8\pm 0.3$ $9.1\pm 0.6$ $13.9\pm 0.8$ $4.3\pm 1.2$
+5. AGEM $4.6\pm 2.8$ $4.5\pm 2.2$ $13.0\pm 0.2$ $2.7\pm 0.9$
+6. RandR $24.9\pm 0.6$ $28.0\pm 0.7$ $10.0\pm 0.5$ $5.7\pm 0.3$
+7. PerpR (Ours) $\underline{26.1}\pm 1.6$ $\underline{28.9}\pm 1.8$ $\underline{9.8}\pm 0.2$ $\textbf{6.1}\pm 0.4$
+8. ESR (Ours) $\textbf{28.2}\pm 0.4$ $\textbf{31.9}\pm 0.9$ $\textbf{9.2}\pm 0.3$ $\underline{5.8}\pm 0.5$
+[…]
+Table 1: Comparative results of various methods on I-CVLN and D-CVLN. This table reports the mean and standard deviation values across three distinct learning curricula.
 5.3 Comparative Results
 The results in Table 1 provide a comparative analysis of various methods on I-CVLN and D-CVLN, highlighting the impact of continual learning (CL) strategies in overcoming catastrophic forgetting during sequential training.
-Performance Bounds: Vanilla and Joint provide performance bounds, where Vanilla reflects a lower bound due to a lack of mechanisms to mitigate forgetting, and Joint represents an upper bound with full access to all data simultaneously.
-The large performance gap between these two methods underscores the difficulty of applying continual learning to CVLN, where each new scene domain introduces unique challenges that increase the risk of forgetting prior knowledge.
+Performance Bounds: Vanilla and Joint define performance bounds, with Vanilla as the lower bound due to no forgetting mitigation, and Joint as the upper bound with full data access.
+Their large performance gap highlights the challenge of continual learning in CVLN, where new scene domains introduce unique challenges and increase forgetting risk.
 Performance of Traditional CL Approaches: Standard CL methods such as L2, A-GEM, and AdapterCL fall short of even the Vanilla baseline, indicating that traditional CL strategies struggle with the complex, dynamic environments of I-CVLN and D-CVLN.
@@
 These findings highlight the adaptability and robustness of PerpR and ESR in handling the unique demands of CVLN tasks.
-Our proposed methods provide a significant improvement in mitigating catastrophic forgetting and adapting to new environments, making them more suitable for complex, multi-domain CVLN tasks than traditional CL methods.
 5.4 Analysis
-5.4.1 Memory Size Analysis
-The proposed methods are rehearsal-based methods that use replay memory.
-Since the performance of these methods can vary depending on the size of the replay memory, we examined the performance differences of rehearsal-based methods according to the size of their replay memory in Table 3.
-PerpR and ESR perform better as the replay memory size increases on both datasets.
-This demonstrates that larger memory can help retain previously learned knowledge by storing a wider variety of data samples.
-5.4.2 Stability-Plasticity Trade-off Analysis
-In CVLN, the agents need to have memory stability and learning plasticity.
-Memory stability ensures retention of previously acquired knowledge, while learning plasticity enables the efficient assimilation of new information.
-These attributes are often in trade-off, known as the stability-plasticity dilemma [33].
-Stability (S) is defined as the agents’ average performance on previously encountered scene domains after training on a new one, while Plasticity (P) represents the average initial performance in newly encountered domains.
-Catastrophic forgetting occurs when plasticity exceeds stability.
-We evaluate the stability-plasticity trade-off [37], the harmonic mean of S and P, visualized in Figure 5.
-ESR achieves a better trade-off in I-CVLN than RandR and Vanilla, with greater plasticity and stability, while PerpR performs similarly in D-CVLN.
-Ablation I-CVLN D-CVLN
+Memory Size Analysis The proposed rehearsal-based methods use replay memory, and their performance depends on its size.
+Table 2 shows that PerpR and ESR perform better with larger memory on both datasets, indicating that larger memory helps retain knowledge by storing more diverse samples.
 […]
-PerpR $\textbf{26.1}\pm 1.6$ $\textbf{28.9}\pm 1.8$ $\textbf{9.8}\pm 0.2$ $\textbf{6.1}\pm 0.4$
-PerpR-Rev. $24.6\pm 1.1$ $27.1\pm 1.2$ $10.2\pm 0.3$ $5.4\pm 0.2$
-Table 4: Impact of memory update reversal on performance in PerpR.
+[…]
+[…]
+PerpR 500 23.8 26.4 9.8 100 6.2
+1000 28.7 31.1 8.7 200 6.0
+[…]
+ESR 500 27.7 31.1 8.8 100 5.5
+1000 29.8 34.4 9.1 200 5.9
+[…]
+Table 2: Comparative analysis of the impact of replay memory size on performance for rehearsal-based methods.
+Larger memory sizes improve performance, indicating a positive correlation between buffer size and agent performance.
+Stability-Plasticity Trade-off Analysis In CVLN, agents require both memory stability to retain past knowledge and learning plasticity to acquire new information.
+These are often in trade-off, known as the stability-plasticity dilemma [Parisi et al.(2019)Parisi, Kemker, Part, Kanan, and Wermter].
+Stability (S) measures average performance on past domains after learning a new one, while Plasticity (P) measures initial performance on new domains.
+Catastrophic forgetting occurs when plasticity outweighs stability.
+We assess this trade-off [Sarfraz et al.(2022)Sarfraz, Arani, and Zonooz] using the harmonic mean of S and P (Figure 4).
+ESR achieves a better trade-off in I-CVLN than RandR and Vanilla, while PerpR shows comparable performance in D-CVLN.
 Figure 4: Stability-plasticity trade-off comparison in I-CVLN and D-CVLN: We calculate stability and plasticity for agents after learning 10 scene domains.
 5.5 Ablation Study
-5.5.1 Reversing PerpR memory update
-Table 4 shows the results of reversing the replay memory update in PerpR.
+Reversing PerpR memory update Table 4 shows the results of reversing the replay memory update in PerpR.
 PerpR is developed based on the assumption that episodes with higher action perplexity would be more beneficial for future learning.
@@
 PerpR’s performance degrades when replay memory updates are reversed by storing low action perplexity episodes and deleting high ones.
-5.5.2 ESR loss ablation
-Ablation I-CVLN D-CVLN
 […]
-$\mathcal{L}_{ESR\_total}$ 27.7 31.1 8.8 5.5
+[…]
+[…]
+PerpR-Rev. $24.6\pm 1.1$ $27.1\pm 1.2$ $10.2\pm 0.3$ $5.4\pm 0.2$
+[…]
+Table 3: Impact of memory update reversal on performance in PerpR.
+[…]
+[…]
+[…]
 - $\mathcal{L}_{M}$ 26.0 27.9 9.7 5.4
 - $\mathcal{L}_{ESR}$ 22.7 24.9 10.4 5.1
-Table 5: Ablation study on ESR for I-CVLN and D-CVLN. This table shows the effects of removing specific loss components on the overall performance.
-Table 5 shows the results of an ablation study on the ESR loss in Equation 7.
+[…]
+Table 4: Ablation study on ESR for I-CVLN and D-CVLN. Effects of removing specific loss components.
+ESR loss ablation Table 4 shows the results of an ablation study on the ESR loss in Equation 7.
 The removal of each of the two losses shows decreased performance across all datasets.
@@
 6 Conclusions and Future Work
-In this paper, we introduced the Continual Vision-and-Language Navigation (CVLN) paradigm to enable VLN agents to learn in changing environments without forgetting previous knowledge.
+In this paper, we introduced the Continual Vision-and-Language Navigation paradigm to enable VLN agents to learn in changing environments without forgetting previous knowledge.
 We proposed two settings, I-CVLN and D-CVLN, to account for various forms of natural language instruction.
@@
 Our experiments showed that these methods outperform existing continual learning approaches, highlighting the importance of effective replay memory in CVLN.
-For future work, we plan to extend CVLN to continuous environments and design new settings and datasets specifically for VLN continual learning, beyond existing dataset-based benchmarks.
+For future work, we plan to extend CVLN to continuous environments and design new settings and datasets specifically for VLN, beyond existing dataset-based benchmarks.
+Acknowledgements This work was partly supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) under grants RS-2021-II212068-AIHub (10%), RS-2021-II211343-GSAI (10%), RS-2022-II220951-LBA (15%), and RS-2022-II220953-PICA (15%); the National Research Foundation of Korea (NRF) under grants RS-2024-00353991-SPARC (15%) and RS-2023-00274280-HEI (15%); the Korea Evaluation Institute of Industrial Technology (KEIT) under grant RS-2024-00423940 (10%); and the Gwangju Metropolitan City (Artificial Intelligence Industrial Convergence Cluster Development Project, 10%) funded by the Korean government.